### v1

@@ 2.2 Variational geometry and differentiation of multifunctions
Let $A$ be a set in $\mathbb{R}^{s}$ and let $\bar{x}\in A$. Then
@@
    The tangent (contingent, Bouligand) cone to $A$ at $\bar{x}$ is given by
    $T_{A}(\bar{x}):=\mathop{{\rm Lim}\,{\rm sup}}\limits_{t\downarrow 0}\frac{A-\bar{x}}{t}$
    and the paratingent cone to $A$ at $\bar{x}$ is given by
    $T^{P}_{A}(\bar{x}):=\mathop{{\rm Lim}\,{\rm sup}}\limits_{{\textstyle{x\mathop{\rightarrow}\limits^{A}\bar{x}\atop t\downarrow 0}}}\frac{A-x}{t}.$
    The regular (Clarke) tangent cone to $A$ at $\bar{x}$ amounts to
    $\widehat{T}_{A}(\bar{x}):=\mathop{{\rm Lim}\,{\rm inf}}_{{\textstyle{x\mathop{\rightarrow}\limits^{A}\bar{x}\atop t\downarrow 0}}}\frac{A-x}{t}.$
@@
    The set
    $\widehat{N}_{A}(\bar{x}):=(T_{A}(\bar{x}))^{\circ},$
    i.e. the (negative) polar cone to the tangent cone $T_{A}(\bar{x})$, is the regular (Fréchet) normal cone to $A$ at $\bar{x}$, and
    $N_{A}(\bar{x}):=\mathop{{\rm Lim}\,{\rm sup}}\limits_{\stackrel{{\scriptstyle A}}{{x\rightarrow\bar{x}}}}\widehat{N}_{A}(x)$
    is the limiting (Mordukhovich) normal cone to $A$ at $\bar{x}$.
If $A$ is convex, then $\widehat{N}_{A}(\bar{x})=N_{A}(\bar{x})$ amounts to the classical normal cone in the sense of convex analysis and we will write $N_{A}(\bar{x})$.
@@
Given a set-valued mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$, we denote by $\mathrm{dom}\,F$ the set of all points $x$ with $F(x)\not=\emptyset$. We call $F$ single-valued, if for every $x\in\mathrm{dom}\,F$ the set $F(x)$ is a singleton. In this case, we can omit the second argument and write $DF(x)$, $\widehat{D}^{*}F(x),\ldots$ instead of $DF(x,F(x))$, $\widehat{D}^{*}F(x,F(x)),\ldots$. Given a mapping $F:\Omega\subset\mathbb{R}^{n}\to\mathbb{R}^{m}$, we can identify it with a set-valued single-valued mapping $\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ via
$x\to\begin{cases}\{F(x)\}&\mbox{if $x\in\Omega$}\\ \emptyset&\mbox{else.}\end{cases}$
However, be aware that when considering limiting objects at $x$ where $F$ is not continuous, it is not enough to consider only sequences $x_{k}\to x$ but we must work with sequences $(x_{k},F(x_{k}))\to(x,F(x))$.
If a mapping $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, is differentiable at $\bar{x}\in\Omega$, there holds
$DF(\bar{x})(u)=\nabla F(\bar{x})u,\ u\in\mathbb{R}^{n}.$
If $F$ is even strictly differentiable at $\bar{x}$, then also
$D_{*}F(\bar{x})(u)=\nabla F(\bar{x})u,\ u\in\mathbb{R}^{n},\quad D^{*}F(\bar{x})(v^{*})=\nabla F(\bar{x})^{T}v^{*},\ v^{*}\in\mathbb{R}^{m}$
holds. A generalization of strict differentiability to general set-valued mappings is provided by the notion of strict proto-differentiability.
Following [27], a set-valued mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is called proto-differentiable at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$ if the set-valued mappings
$\Delta_{\bar{x},\bar{y},t}F:\xi\mapsto\big{(}F(\bar{x}+t\xi)-\bar{y})/t\quad\mbox{ for $t>0$}$
graph-converge as $t\downarrow 0$, i.e., their graphs converge as subsets of $\mathbb{R}^{n}\times\mathbb{R}^{m}$. Of course, the limit must be equal to the graphical derivative $DF(\bar{x},\bar{y})$. This is the same as requiring that
$\mathop{{\rm Lim}\,{\rm sup}}_{t\downarrow 0}\frac{\mathrm{gph}\,F-(\bar{x},\bar{y})}{t}=\mathop{{\rm Lim}\,{\rm inf}}_{t\downarrow 0}\frac{\mathrm{gph}\,F-(\bar{x},\bar{y})}{t}\big{(}=\mathrm{gph}\,DF(\bar{x},\bar{y})\big{)}.$
According to [25] we say that $F$ is strictly proto-differentiable at $(\bar{x},\bar{y})$ if, more generally, the mappings
@@
graph-converge as $t\downarrow 0$ and $(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})$, the limit being $D_{*}F(\bar{x},\bar{y})$. This means that,
$\mathop{{\rm Lim}\,{\rm sup}}_{{\textstyle{t\downarrow 0\atop(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}}}\frac{\mathrm{gph}\,F-(x,y)}{t}=\mathop{{\rm Lim}\,{\rm inf}}_{{\textstyle{t\downarrow 0\atop(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}}}\frac{\mathrm{gph}\,F-(x,y)}{t}\big{(}=\mathrm{gph}\,D_{*}F(\bar{x},\bar{y})\big{)}.$ (3)
@@
If $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is strictly proto-differentiable at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$ then $T_{\mathrm{gph}\,F}(\bar{x},\bar{y})$ is a subspace of $\mathbb{R}^{n}\times\mathbb{R}^{m}$ and
$\mathrm{gph}\,DF(\bar{x},\bar{y})$ $=T_{\mathrm{gph}\,F}(\bar{x},\bar{y})=\mathop{{\rm Lim}\,{\rm sup}}_{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}T_{\mathrm{gph}\,F}(x,y)=\mathop{{\rm Lim}\,{\rm inf}}_{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}T_{\mathrm{gph}\,F}(x,y)=\widehat{T}_{\mathrm{gph}\,F}(\bar{x},\bar{y})$ (4)
$=\mathrm{gph}\,D_{*}F(\bar{x},\bar{y}).$
@@
By taking into account [28, Theorem 6.26], we obtain from (3) that
$\mathrm{gph}\,DF(\bar{x},\bar{y})$ $=T_{\mathrm{gph}\,F}(\bar{x},\bar{y})\subset\mathrm{gph}\,D_{*}F(\bar{x},\bar{y})=\mathop{{\rm Lim}\,{\rm sup}}_{{\textstyle{t\downarrow 0\atop(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}}}\frac{\mathrm{gph}\,F-(x,y)}{t}$
$=\mathop{{\rm Lim}\,{\rm inf}}_{{\textstyle{t\downarrow 0\atop(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}}}\frac{\mathrm{gph}\,F-(x,y)}{t}=\mathop{{\rm Lim}\,{\rm inf}}_{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}T_{\mathrm{gph}\,F}(x,y)=\widehat{T}_{\mathrm{gph}\,F}(\bar{x},\bar{y})\subset T_{\mathrm{gph}\,F}(\bar{x},\bar{y}).$
Hence, in order to prove (4) there remains to show
$A:=\mathop{{\rm Lim}\,{\rm sup}}_{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}T_{\mathrm{gph}\,F}(x,y)=\mathop{{\rm Lim}\,{\rm sup}}_{{\textstyle{t\downarrow 0\atop(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}}}\big{(}\mathrm{gph}\,F-(x,y)\big{)}/t=:B.$
Clearly, by the inclusions above we have $B=T_{\mathrm{gph}}\,F(\bar{x},\bar{y})\subset A$. Now consider $(u,v)\in A$ together with sequences $(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})$ and $(u_{k},v_{k})\in T_{\mathrm{gph}\,F}(x_{k},y_{k})$ satisfying $(u_{k},v_{k})\to(u,v)$. By definition of the tangent cone, for every $k$ we can find some $t_{k}\in(0,\frac{1}{k})$ and some $(\hat{x}_{k},\hat{y}_{k})\in\mathrm{gph}\,F$ with
$\|\frac{(\hat{x}_{k},\hat{y}_{k})-(x_{k},y_{k})}{t_{k}}-(u_{k},v_{k})\|\leq\frac{1}{k}$
implying $\big{(}(\hat{x}_{k},\hat{y}_{k})-(x_{k},y_{k})\big{)}/t_{k}\to(u,v)$. This proves $(u,v)\in B$ and the equality $A=B$ is established.
Finally, in order to show that $T_{\mathrm{gph}\,F}(\bar{x},\bar{y})$ is a subspace, observe first that $T_{\mathrm{gph}\,F}(\bar{x},\bar{y})=B=\widehat{T}_{\mathrm{gph}\,F}(\bar{x},\bar{y})$ is a convex cone. Further, for any $(u,v)\in B=T_{\mathrm{gph}\,F}(\bar{x},\bar{y})$ there are sequences $(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})$, $t_{k}\downarrow 0$ and $(\hat{x}_{k},\hat{y}_{k})\in T_{\mathrm{gph}\,F}(x_{k},y_{k})$ with $\big{(}(\hat{x}_{k},\hat{y}_{k})-(x_{k},y_{k})\big{)}/t_{k}\to(u,v)$. Then there also holds $(\hat{x}_{k},\hat{y}_{k}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})$ and
$-(u,v)=\lim_{k\to\infty}\frac{(x_{k},y_{k})-(\hat{x}_{k},\hat{y}_{k})}{t_{k}}\in B$
follows. Hence $B=-B$ and it follows that $B$ is a subspace. ∎
In case of single-valued Lipschitzian maps, strict proto-differentiability coincides with strict differentiability.
@@
exist and equal to $D_{*}F(\bar{x})(u)$. Hence, if $F$ is strictly proto-differentiable at $(\bar{x},F(\bar{x}))$ then $D_{*}F(\bar{x})$ is single-valued, $\mathrm{dom}\,D_{*}F(\bar{x})=\mathbb{R}^{n}$ and $\mathrm{gph}\,D_{*}F(\bar{x})$ is a subspace by Lemma 2.4. We conclude that $D_{*}F(\bar{x},\bar{y})$ is in fact a linear mapping from $\mathbb{R}^{n}$ to $\mathbb{R}^{m}$ and therefore there exists an $m\times n$ matrix $A$ with $D_{*}F(\bar{x})(u)=Au$, $u\in\mathbb{R}^{n}$. Thus $F$ is strictly differentiable at $\bar{x}$. The reverse implication easily follows from the definition of strict differentiability and the fact that $D_{*}\nabla F(\bar{x})(u)=\nabla F(\bar{x})u$, $u\in\mathbb{R}^{n}$, in this case. ∎
@@ 2.3 Closure of mappings
It is well-known that the Clarke subdifferential $\partial_{c}\theta$ of a Lipschitz continuous function $\theta:\Omega\to\mathbb{R}$, $\Omega\subset\mathbb{R}^{n}$ open, has the following properties:
  • $\partial_{c}\theta$ is locally bounded
  • The set $\partial_{c}\theta(x)$ is nonempty, convex and compact for every $x\in\Omega$.
  • $\partial_{c}\theta$ is upper semicontinuous on $\Omega$.
We now discuss a possibility how to extend an arbitrary locally bounded mapping to a mapping having the properties stated above.
@@
In most cases we consider set-valued mappings between Euclidean spaces.
Definition 2.6.
Consider a mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$.
  1. 1.
    $\Psi$ is called locally bounded at $\bar{x}\in\mathbb{R}^{n}$ if there is a neighborhood $U$ of $\bar{x}$ and a nonnegative real $M$ such that
    $\|y\|\leq M\ \forall y\in\Psi(x)\ \forall x\in U.$ (5)
    We denote by $\mathrm{bnd}\,\Psi(\bar{x})$ the infimum of all reals $M$ such that (5) holds for some neighborhood $U$.
  2. 2.
    $\Psi$ is called outer semicontinuous (osc) at $\bar{x}\in\mathbb{R}^{n}$ if $(\mathrm{cl}\,\Psi)(\bar{x})=\Psi(\bar{x})$.
  3. 3.
    $\Psi$ is called upper semicontinuous (usc) at $\bar{x}$ if $\Psi(\bar{x})$ is closed and for every open set $V\supset\Psi(\bar{x})$ the set $\{x\,\big{|}\,\Psi(x)\subset V\}$ is a neighborhood of $\bar{x}$.
These terms are invoked on $X$, a subset of $\mathbb{R}^{n}$, when the property holds for every $x\in X$. The set $X$ will not be mentioned if $X=\mathbb{R}^{n}$.
By the definition, the function $x\mapsto\mathrm{bnd}\,\Psi(x)$ is usc.
Given $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$, we denote by $\rm cocl\,\Psi$ the mapping defined by
@@
It is easy to see that $\mathrm{dom}\,(\rm cocl\,\Psi)=\mathrm{dom}\,(\mathrm{cl}\,\Psi)$.
Lemma 2.7.
Let $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ be a mapping. If $\Psi$ is locally bounded at $x\in\mathbb{R}^{n}$ then $\mathrm{cl}\,\Psi$ and $\rm cocl\,\Psi$ are also locally bounded at $x$ and $(\rm cocl\,\Psi)(x)$ is a convex and compact set. Further, both $\mathrm{cl}\,\Psi$ and $\rm cocl\,\Psi$ are both osc and usc at $x$.
Proof.
Let $\Psi$ be locally bounded at $x\in\mathbb{R}^{n}$ and consider an open neighborhood $U$ of $x$ such that $V:=\bigcup_{x^{\prime}\in U}\Psi(x^{\prime})$ is bounded. We claim that
$\bigcup_{x^{\prime}\in U}(\mathrm{cl}\,{\Psi})(x^{\prime})\subset\bigcup_{x^{\prime}\in U}(\rm cocl\,{\Psi})(x^{\prime})\subset\mathrm{conv}\,(\mathrm{cl}\,V).$
Indeed, consider a point $(x^{\prime},y^{\prime})\in\mathrm{gph}\,(\mathrm{cl}\,\Psi)\cap(U\times\mathbb{R}^{m})$ together with a sequence $(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,\Psi}}(x^{\prime},y^{\prime})$. Then for all $k$ sufficiently large we have $x_{k}\in U$ and consequently $y_{k}\in V$. Thus $y^{\prime}\in\mathrm{cl}\,V$ and $\bigcup_{x^{\prime}\in U}(\rm cocl\,{\Psi})(x^{\prime})\subset\mathrm{conv}\,(\mathrm{cl}\,V)$ follows. This verifies the local boundedness of $\rm cocl\,\Psi$ at $x$. Since the set $(\mathrm{cl}\,\Psi)(x)$ is closed and bounded, it is compact and the same holds true for its convex hull $(\rm cocl\,\Psi)(x)$.
Due to the local boundedness of $\mathrm{cl}\,\Psi$ and $\rm cocl\,\Psi$ at $x$, the mappings $\mathrm{cl}\,\Psi$ and $\rm cocl\,\Psi$ are osc if and only if they are usc, cf. [28, Theorem 5.19]. Clearly, $\mathrm{cl}\,\Psi$ is osc at $x$ and there remains to show that $\rm cocl\,\Psi$ is usc at $x$. If $x\not\in\mathrm{dom}\,(\rm cocl\,\Psi)=\mathrm{dom}\,(\mathrm{cl}\,\Psi)$ then the set $\{x^{\prime}\,\big{|}\,x^{\prime}\not\in\mathrm{dom}\,(\mathrm{cl}\,\Psi)=\mathrm{dom}\,(\rm cocl\,\Psi)\}$ is a neighborhood of $x$ since $\mathrm{cl}\,\Psi$ is usc at $x$. Now let $x\in\mathrm{dom}\,(\rm cocl\,\Psi)$ and consider an open set $V\supset(\rm cocl\,{\Psi})(x)$. Since $(\rm cocl\,{\Psi})(x)$ is compact, we can find some $\epsilon>0$ such that the convex open set $(\rm cocl\,{\Psi})(x)+\epsilon{\cal B}$ is contained in $V$. Since the mapping $\mathrm{cl}\,\Psi$ is usc, the set $U:=\{x^{\prime}\,\big{|}\,(\mathrm{cl}\,\Psi)(x^{\prime})\subset(\mathrm{cl}\,\Psi)(x)+\epsilon{\cal B}\}$ is a neighborhood of $x$. For every $x^{\prime}\in U$ we have
$(\rm cocl\,{\Psi})(x^{\prime})=\mathrm{conv}\,\big{(}(\mathrm{cl}\,\Psi)(x^{\prime})\big{)}\subset\mathrm{conv}\,\big{(}(\mathrm{cl}\,\Psi)(x)+\epsilon{\cal B}\big{)}=(\rm cocl\,{\Psi})(x)+\epsilon{\cal B}\subset V$
and the lemma is proved. ∎
@@
Let $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, be locally Lipschitzian and consider the mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m\times n}$ given by
$\Psi_{F}(x):=\begin{cases}\{\nabla F(x)\}&\mbox{if $x\in{\cal O}_{F}$}\\ \emptyset&\mbox{else.}\end{cases}$
Then for every $x\in\Omega$ there holds ¯$\overline{\nabla}F(x)=(\mathrm{cl}\,\Psi_{F})(x)$ and $\mathrm{conv}\,\overline{\nabla}F(x)=(\rm cocl\,\Psi_{F})(x)$.

### v2

@@ 2.2 Variational geometry and differentiation of multifunctions
Let $\Omega\subset\mathbb{R}^{s}$ and $\bar{x}\in\Omega$.
@@
    The tangent (contingent, Bouligand) cone $T_{\Omega}(\bar{x})$, the paratingent cone $T^{P}_{\Omega}(\bar{x})$ and the regular (Clarke) tangent cone $\widehat{T}_{\Omega}(\bar{x})$ to $\Omega$ at $\bar{x}$ are given by
    $T_{\Omega}(\bar{x}):=\mathop{{\rm Lim}\,{\rm sup}}\limits_{t\downarrow 0}\frac{\Omega-\bar{x}}{t},\quad T^{P}_{\Omega}(\bar{x}):=\mathop{{\rm Lim}\,{\rm sup}}\limits_{{\textstyle{x\mathop{\rightarrow}\limits^{\Omega}\bar{x}\atop t\downarrow 0}}}\frac{\Omega-x}{t},\quad\widehat{T}_{\Omega}(\bar{x}):=\mathop{{\rm Lim}\,{\rm inf}}_{{\textstyle{x\mathop{\rightarrow}\limits^{\Omega}\bar{x}\atop t\downarrow 0}}}\frac{\Omega-x}{t}.$
@@
    The set $\widehat{N}_{\Omega}(\bar{x}):=(T_{\Omega}(\bar{x}))^{\circ}$, i.e. the (negative) polar cone to the tangent cone $T_{\Omega}(\bar{x})$, is the regular (Fréchet) normal cone to $\Omega$ at $\bar{x}$, and
    $N_{\Omega}(\bar{x}):=\mathop{{\rm Lim}\,{\rm sup}}\limits_{x\mathop{\rightarrow}\limits^{\Omega}\bar{x}}\widehat{N}_{\Omega}(x)$
    is the limiting (Mordukhovich) normal cone to $\Omega$ at $\bar{x}$.
If $\Omega$ is convex, then $\widehat{N}_{\Omega}(\bar{x})=N_{\Omega}(\bar{x})$ amounts to the classical normal cone in the sense of convex analysis and we will write $N_{\Omega}(\bar{x})$.
@@
We call $F$ single-valued, if for every $x\in\mathrm{dom}\,F$ the set $F(x)$ is a singleton. In this case, we can omit the second argument and write $DF(x)$, $\widehat{D}^{*}F(x),\ldots$ instead of $DF(x,F(x))$, $\widehat{D}^{*}F(x,F(x)),\ldots$.
If a mapping $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, is differentiable at $\bar{x}\in\Omega$ then $DF(\bar{x})(u)=\nabla F(\bar{x})u,$ $u\in\mathbb{R}^{n}.$ If $F$ is even strictly differentiable at $\bar{x}$, then also
$D_{*}F(\bar{x})(u)=\nabla F(\bar{x})u,\ u\in\mathbb{R}^{n},\quad D^{*}F(\bar{x})(v^{*})=\nabla F(\bar{x})^{T}v^{*},\ v^{*}\in\mathbb{R}^{m}$ (2.4)
holds, cf. [35]. A generalization of strict differentiability to general set-valued mappings is provided by the notion of strict proto-differentiability. According to [32] we say that a set-valued mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is strictly proto-differentiable at $(\bar{x},\bar{y})$ if the mappings
@@
graph-converge as $t\downarrow 0$ and $(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})$, the limit being $D_{*}F(\bar{x},\bar{y})$. This means that
$\mathop{{\rm Lim}\,{\rm sup}}_{{\textstyle{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})\atop t\downarrow 0}}}\frac{\mathrm{gph}\,F-(x,y)}{t}=\mathop{{\rm Lim}\,{\rm inf}}_{{{\textstyle{(\atop x}},y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}{t\downarrow 0}}\frac{\mathrm{gph}\,F-(x,y)}{t}\big{(}=\mathrm{gph}\,D_{*}F(\bar{x},\bar{y})\big{)}.$ (2.5)
@@
If $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is strictly proto-differentiable at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$ and $\mathrm{gph}\,F$ is locally closed at $(\bar{x},\bar{y})$ then $T_{\mathrm{gph}\,F}(\bar{x},\bar{y})$ is a subspace of $\mathbb{R}^{n}\times\mathbb{R}^{m}$ and
$\mathrm{gph}\,DF(\bar{x},\bar{y})$ $=T_{\mathrm{gph}\,F}(\bar{x},\bar{y})=\mathop{{\rm Lim}\,{\rm sup}}_{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}T_{\mathrm{gph}\,F}(x,y)=\mathop{{\rm Lim}\,{\rm inf}}_{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}T_{\mathrm{gph}\,F}(x,y)$ (2.6)
$=\widehat{T}_{\mathrm{gph}\,F}(\bar{x},\bar{y})=\mathrm{gph}\,D_{*}F(\bar{x},\bar{y}).$
@@
By taking into account [35, Theorem 6.26], we obtain from (2.5) that
$\mathrm{gph}\,DF$ $(\bar{x},\bar{y})=T_{\mathrm{gph}\,F}(\bar{x},\bar{y})\subset\mathrm{gph}\,D_{*}F(\bar{x},\bar{y})=\mathop{{\rm Lim}\,{\rm sup}}_{{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})},\ {t\downarrow 0}}\frac{\mathrm{gph}\,F-(x,y)}{t}$
$=\mathop{{\rm Lim}\,{\rm inf}}_{{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})},\ {t\downarrow 0}}\frac{\mathrm{gph}\,F-(x,y)}{t}=\mathop{{\rm Lim}\,{\rm inf}}_{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}T_{\mathrm{gph}\,F}(x,y)=\widehat{T}_{\mathrm{gph}\,F}(\bar{x},\bar{y})\subset T_{\mathrm{gph}\,F}(\bar{x},\bar{y}).$
Hence, in order to prove (2.6) there remains to show the equality
$A:=\mathop{{\rm Lim}\,{\rm sup}}_{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}T_{\mathrm{gph}\,F}(x,y)=\mathop{{\rm Lim}\,{\rm sup}}_{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y}),\ t\downarrow 0}\big{(}\mathrm{gph}\,F-(x,y)\big{)}/t=:B.$
Clearly, by the inclusions above we have $B=\mathrm{gph}\,D_{*}F(\bar{x},\bar{y})=T_{\mathrm{gph}}\,F(\bar{x},\bar{y})\subset A$. By [18, Proposition 2.1(b)] the reverse inclusion $A\subset B$ always holds and the equality $A=B$ is established.
Finally, in order to show that $T_{\mathrm{gph}\,F}(\bar{x},\bar{y})$ is a subspace, observe first that $T_{\mathrm{gph}\,F}(\bar{x},\bar{y})=\mathrm{gph}\,D_{*}F(\bar{x},\bar{y})=\widehat{T}_{\mathrm{gph}\,F}(\bar{x},\bar{y})$ is a convex cone. By [18, Proposition 2.1(a)], the strict derivative also satisfies $\mathrm{gph}\,D_{*}F(\bar{x},\bar{y})=-\mathrm{gph}\,D_{*}F(\bar{x},\bar{y})$ and thus $T_{\mathrm{gph}\,F}(\bar{x},\bar{y})$ is a subspace. ∎
In case of single-valued Lipschitz maps, strict proto-differentiability coincides with strict differentiability.
@@
exist and are equal to $D_{*}F(\bar{x})(u)$. Hence, if $F$ is strictly proto-differentiable at $(\bar{x},F(\bar{x}))$ then $D_{*}F(\bar{x})$ is single-valued, $\mathrm{dom}\,D_{*}F(\bar{x})=\mathbb{R}^{n}$ and $\mathrm{gph}\,D_{*}F(\bar{x})$ is a subspace by Lemma 2.4. We conclude that $D_{*}F(\bar{x},\bar{y})$ is in fact a linear mapping from $\mathbb{R}^{n}$ to $\mathbb{R}^{m}$ and therefore there exists an $m\times n$ matrix $A$ with $D_{*}F(\bar{x})(u)=Au$, $u\in\mathbb{R}^{n}$. Thus $F$ is strictly differentiable at $\bar{x}$. The reverse implication easily follows from the definition of strict differentiability and (2.4). ∎
@@ 2.3 Closure of mappings
It is well-known that the Clarke subdifferential $\partial^{c}\theta$ of a Lipschitz continuous function $\theta:\Omega\to\mathbb{R}$, $\Omega\subset\mathbb{R}^{n}$ open, is upper semicontinuous (usc) and has nonempty, convex and compact values $\partial^{c}\theta(x)$ for every $x\in\Omega$. We now discuss a possibility how to extend an arbitrary locally bounded mapping to a mapping having these properties.
@@
In most cases we consider set-valued mappings between Euclidean spaces. The mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is called locally bounded at $\bar{x}\in\mathbb{R}^{n}$ if there is a neighborhood $U$ of $\bar{x}$ and a nonnegative real $M$ such that
$\|y\|\leq M\ \forall y\in\Psi(x)\ \forall x\in U.$ (2.7)
We denote by $\mathrm{bnd}\,\Psi(\bar{x})$ the infimum of all reals $M$ such that (2.7) holds for some neighborhood $U$. By the definition, the function $x\mapsto\mathrm{bnd}\,\Psi(x)$ is usc. Further, we denote by ${\rm cocl}\,\Psi$ the mapping defined by
@@
It is easy to see that $\mathrm{dom}\,({\rm cocl}\,\Psi)=\mathrm{dom}\,(\mathrm{cl}\,\Psi)$. Moreover, if $\Psi$ is locally bounded at $x\in\mathbb{R}^{n}$ then $\mathrm{cl}\,\Psi$ and ${\rm cocl}\,\Psi$ are also locally bounded at $x$ and $({\rm cocl}\,\Psi)(x)$ is a convex and compact set. Further, both $\mathrm{cl}\,\Psi$ and ${\rm cocl}\,\Psi$ are usc at $x$.
@@
Let $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, be locally Lipschitz and consider the mapping $\Psi_{F}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m\times n}$ given by $\Psi_{F}(x):=\{\nabla F(x)\},\ x\in\mathrm{dom}\,\Psi_{F}:={\cal O}_{F}$. Then for every $x\in\Omega$ there holds $(\mathrm{cl}\,\Psi_{F})(x)=\overline{\nabla}F(x)$ and $({\rm cocl}\,\Psi_{F})(x)=\mathrm{conv}\,\overline{\nabla}F(x)$.